Applied vs. Actual Manufacturing Overhead Sloan Manufacturing Corporation applies manufacturing overhead on the basis of 120% of direct labor cost. An analysis of the related accounts and job order cost sheet indicates that during the year total manufacturing overhead incurred was $525,000 and that at year-end Work in Process Inventory, Finished Goods Inventory, and Cost of Goods Sold included $75,000, $50,000, and $375,000, respectively, of direct labor incurred during the current year a. Determine the over-applied manufacturing overhead at year-end (assume it is significant). Applied Manufacturing Overhead Work in process Cost of goods sold Total: Over-applied manufacturing overhead s b. Prepare a journal entry to record the disposition of the over-applied manufacturing overhead General Journal Description Debit Credit
Answer #1

a.Manufacturing Overhead is applied at 120% of direct labor cost.

Direct Labor cost of Work in process =$75000.

So Applied manufacturing overhead for work in process will be 120% of $75000 = $90,000

Direct Labor cost of Finished Goods = $50000

So Applied manufacturing overhead for Finished Goods will be 120% of $50000 = $60000.

Direct Labor cost of Cost of goods sold = $375000.

So Applied manufacturing overhead for cost of goods sold will be 120% of $375000 = $450,000

Total Applied manufacturing overhead is $90000 + $60000+ $450000= $600,000

Total manufacturing overhead incurred is $525000.

Over-applied manufacturing overhead is $600000- $525000 = $75000

b. When the manufacturing overhead is over applied the account will have a credit balance. To dispose the over applied amount, the manufacturing overhead account will be debited with the over applied amount and the credit will be to the 3 accounts as per their proportions.

The account balances after applying manufacturing over head are

Work in process - $75000 + $90000 = $165,000

Finished Goods - $50000+ $60000 = $110000

Cost of Goods Sold - $375000 + $450000= $825,000

Total of 3 accounts is $165000 + $110000 + $825000 = $1,100,000

Prorata of these 3 accounts will be

Work in process $165000/$1100000= 0.15

Finished Goods $110000/$1,100,000 = 0.10

Cost of Goods Sold $825000/$1,100,000 = 0.75

The over applied amount will be credited to these 3 accounts as below

Work in process $75,000 * 0.15 = $11,250

Finished Goods $75,000 * 0.10 = $7500

Cost of Goods Sold $75,000 * 0.75 = $56,250

The Journal entry will be

Manufacturing Overhead A/c     Dr.                         $75000

To Work in process A/c                  Cr.                                          $11250

To Finished Goods A/c                   Cr.                                          $7500

To Cost of Goods Sold A/c            Cr.                                          $56250

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU